OpenStudy (anonymous):

HELP!! Sabrina is fencing her yard, which is shaped as a rhombus. She used (6x – 7) meters of fencing material for one side and (3x+ 5) meters for the adjacent side. What is the perimeter of the yard?

OpenStudy (anonymous):

Remember, in a rhombus, the sides are equal. So does this give you a hint?

OpenStudy (anonymous):

Another hint: One side is (6x - 7) and another is (3x + 5). Can you equate these expressions and find x?

Parth (parthkohli):

Hey Pratu! You are right! Although, as the question tells to find the perimeter, we must use the formula. \(\Large \color{MidnightBlue}{\Rightarrow 4(side) = Perimeter(Rhombus) }\) We can use any one of those expressions.

OpenStudy (anonymous):

I know, if we find x, we will be able to find the perimeter as a constant and NOT in terms of x.

Parth (parthkohli):

Well, that is correct, indeed.

Parth (parthkohli):

And then, not to forget, you must substitute x in any one of the expressions.

OpenStudy (anonymous):

The formula to determine the perimeter of a rhombus is 4s and each side (s) of a rhombus is of equal length. So that means: (6x-7)=(3x+5) 6x=3x+12 3x=12 x=4 Substitute 4 for x: 6(4)-7=17 3(4)+5=17 So each side is 17 units: 4s 4(17)= 68 units = the perimeter of the rhombus